Consider a no-load mutual fund with $200 million in assets and 10 million shares at the start of the year and with $250 million in assets and 11 million shares at the end of the year. During the year investors have received income distributions of $2 per share and capital gain distributions of $0.55 per share. Assuming that the fund carries no debt, and that the total expense ratio is 1%, what is the rate of return on the fund
Answer #1

NAV = Fund Asset - Fund Liabilities Outstanding Shares

NAV0 = 200/10 = 20
NAV1 = [250 - (250 × 0.01)]/11 = 22.50

Gross return = (22.50 - 20 + 2 + 0.55) / 20

= 25.25%